The stock of Townships Ski Resorts Inc. just paid a dividend of $0.78. What is the expected capital gains yield if the stock is selling for $28.25 today and the required rate of return is 15 percent?
1
Expert's answer
2019-11-11T15:30:00-0500

Expected capital gains yield formula is CGY = (P1 - P0)/P0,

P0 = D/r = 0.78/0.15 = 5.2,

CGY = (28.25 - 5.2)/5.2 = 4.43 or 443%.